OpenStudy (anonymous):

f(x) = x + 8 and g(x) = x^2 – 6x – 7. Find f(g(2))

OpenStudy (mertsj):

f(g(x))= x^2-6x-7 +8=x^2-6x+1 f(g(2))=(2)^2-6(2)+1= -7

OpenStudy (mertsj):

Alternatively, you could find the g(2) which is (2)^2 -6(2) -7= 4-12-7=-15 and then put that value into the f(x) and get -15+8=-7

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw